OpenStudy (karinamtz0213):

In the given equation, for any real number x and y, what is the value of a+b? -(y-x)+ 2(3x+7)=by+ax+14 a. 5 b. 6 c. 7 d. 8 e. 9

OpenStudy (anonymous):

We have : \[-(y-x)+2(3x+7)=-y+x+6x+14=-y+7x+14\] hence : \[b=-1\\a=7\]

OpenStudy (karinamtz0213):

So would it be letter b: 6?

OpenStudy (anonymous):

What do you think ?! Of course It is b ! :P

OpenStudy (karinamtz0213):

okay thanks

OpenStudy (anonymous):

You are welcome !